OpenStudy (anonymous):

solve for x and y by substitution: x/3+y/4=2 2x/3-y/2=0

OpenStudy (anonymous):

4x +3y= 24 4x -3y= 0 adding we get 8x=24 x=3 put x in first equation 12 +3y=24 3y=12 y=4

OpenStudy (anonymous):

did u like the answer?

OpenStudy (anonymous):

yes thank you sosososo much- im a fan of yours now too! =} lol

OpenStudy (anonymous):

but wait, whered you get he 24 from?

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i replied in your other question

OpenStudy (anonymous):

alrighty